OpenStudy (anonymous):

can someone explain: Simplify the expression. (8 + 7a) + 4? i keep getting 19a. what am i doing wrong?

OpenStudy (lgbasallote):

you added 8 + 7 + 4

OpenStudy (lgbasallote):

you can only add 8 and 14..you cannot touch 7a <---because it has an a

OpenStudy (lgbasallote):

i mean add 8 and 4 not 14

OpenStudy (anonymous):

so whenever a letter is involved i don't touch it because it is a variable?

OpenStudy (cwtan):

For example 1+2+3(4)=is it 6(4)? no it is=3+3(4)

OpenStudy (lgbasallote):

yes @8887

OpenStudy (lgbasallote):

unless another term with the EXACT same letter exists..then you can combine those

OpenStudy (cwtan):

For this i give another example 1(2)+3(2)=4(2)